OpenStudy (anonymous):

Need help with this math question, can someone please help me?. If f(x)=2x^2-x-6 and g(x)=x^2-4, find f(x)/g(x) Answers: 2x+3/x-2 2x-3/x+2 2x+3/x+2 2x-3/x-2

OpenStudy (er.mohd.amir):

f(x)=2x^2-x-6=2x*x-4x+3x-6=2x(x-2)+3(x-2)=(x-2)(2x+3) and x*x-4=(x+2)(x-2) divide it

OpenStudy (anonymous):

okay mmm how would i do that? Im sorry math is my worst subject @Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

what is this factor

OpenStudy (anonymous):

idk /: @Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

(2x-3)(x-2) / (x-2)(x+2) =(2x-3)/(x+2)

OpenStudy (anonymous):

okay thsnk you so much!(:

OpenStudy (anonymous):

@Er.Mohd.AMIR